OpenStudy (anonymous):

perform the operation and simplify: (3+2i/2-3i)

OpenStudy (dumbcow):

\[\frac{(3+2i)}{(2-3i)}*\frac{(2+3i)}{(2+3i)}\] this will eliminate "i" from denominator remember i^2 = -1

OpenStudy (anonymous):

I don't understand how to factor this out because I thought I had to multiple 3i by both 2 and the other 3i

OpenStudy (anonymous):

my teacher said the answer is I. I just don't know how to find it

OpenStudy (dumbcow):

just distribute \[\rightarrow \frac{6+13i+6i^{2}}{4-9i^{2}}\]

OpenStudy (dumbcow):

\[=\frac{13i}{13} = i\]

OpenStudy (anonymous):

how did 4-9i^2 go to 13?

OpenStudy (dumbcow):

i^2 = -1 4-(-9) = 4+9 = 13

OpenStudy (anonymous):

oh yeah you just told me that. thank you!

OpenStudy (dumbcow):

yw